THE EYES OF CERTAIN MEMBER OF A REPTILE FAMILY PASS A VISUAL SIGNAL TO THE BRAIN,WHEN THE VISUAL RECEPTORS ARE STRUCK BY PHOTONS OF WAVELENGHT 850NM.IF A TOTAL ENERGY OF 3.15*10^-14 IS REQUIRED TO TRIP THE SIGNAL, WHAT IS THE MINIMUM NUMBER OF PHOTONS THAT MUST STRIKE TE RECEPTORS?

Grade:11

1 Answers

Arun
5 years ago
E = hc/λ = (6.626×10^-34)*(2.998×10^8) / (890×10^-9) = 2.23×10^-19 J 
Number of photons = Total energy / Energy per photon 
N = 3.15×10^-14 / 2.23×10^-19 = 1.41×10^5 photons (approximately)